OpenStudy (anonymous):

Two dice are thrown simultaneously. Given that sum of the numbers is NOT more than 5, what is the probability that sum is more than 3? 7 10 3 10 5 6 3 4 9 10

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

n(S)= 36 Let A denotes the event of getting even number S THE SUM Ii.e, 2,4,6,8,10 12 Hence A={( 1,1 ),(1,3 ) (1,5 ) (2,2 ) , (2,4 ) , (2,6 ), ( 4,4 ), (4,6 ), ( 4,2 ), (5,3 ),( 3,5 ),( 6,6 ), (6,4 ),.( 5,1 ), (3,1 ) ,( 6,2 ),( 3,3 ), (5,5 ) }, n(A)= 18 thus,required probability= P(A)= n(A)/n(S)= 18/36= 1/2 =50%

OpenStudy (anonymous):

Hope you understand all dat